Identify the focus and directrix of the parabol:
1.) 3x^2= -y

Answer by lwsshak3(11628)

Identify the focus and directrix of the parabola:
1.) 3x^2= -y
..
3x^2= -y
x^2=-y/3
This is of the standard form: x^2=-4py
for given equation
4p=1/3
p=1/12
Given equation is a parabola that opens downward with axis of symmetry, x=0, and vertex at (0,0)
The directrix is a line, y=1/12
The focus is at (0,-1/12)